# Ultra-stretchable and Self-Healable Vitrimers with Tuneable Damping and Mechanical Response

### v1

@@ II.6 Mechanical damping
The -dependent loss tangent is plotted for all semicured (dashed lines) and fully cured (solid lines) vitrimer samples in Figure 5(h). A maximum loss factor () with a value between 2 and 3 is consistently observed. As the cross-link density increases upon full curing, the -relaxation slows down, as shown by the shift of the (Fig. 5g) and (Fig. 5h) loss peaks to higher temperatures, which is consistent with the increase in (see Table 1). Also, increasing the cross-link density, either by addition of cross-linker during synthesis or by curing, shifts the peak in the loss tangent, and both curing and nominal cross-linking concentration can thus be used to control the -position of the mechanical loss maximum. The temperature span of efficient mechanical loss ( 0.3) cover 35 5 °C, which is comparable to the ranges (°C) for commercial styrene-butadiene and natural rubbers ( 2.0). [72, 73] We note that the temperature range for efficient mechanical loss could be further increased by blending of semi-cured and fully cured vitrimers, thus producing materials with efficient loss factors across a range of 35 to 55 °C; such temperature ranges are comparable to those achieved for copolymer-based elastomers ( 1.7), [74] however, without the inherent compatibility issues that exist for blends of different rubbers ( 2.0). [75]
@@
II.7 Room temperature self-healing
@@
The photos in Figure 6(c) show a representative VMA-1’ vitrimer sample before deformation at 0% strain (left), after 1 hour at a strain of 2500 (centre), and after 24 hours at a strain of 3000 (right). In the left and centre photos, the cut can still be observed, but after 24 hours of room temperature self-healing it is no longer easily distinguishable, as shown in the right photo. Thus, we conclude that our semi-cured vitrimers demonstrate good self-healing properties at room temperature, without the addition of water or any other ‘solvent’. We do note that due to the slower segmental dynamics (ca. 20 °C higher ) and slower (circa two orders of magnitude; see Figure S8) network rearrangments, the fully cured samples do not exhibit noticeable room temperature self-healing on a time-scale of 24 hours. However, fully cured vitrimers have the potential to self-heal at temperatures above room temperature.
@@ III Conclusion
We here synthesize and characterise a new vitrimer system based on poly(methyl acrylate) with dynamic crosslinks utilising the associative exchange reactions of dioxaboralane metathesis. We demonstrate that these vitrimers combine ultra-stretchability and mechanical toughness with high mechanical dissipation and room temperature self-healing. Vitrimers with nominal cross-link densities between 2 and 8 mol% were prepared according to two different preparation routes: the first route involves a single curing stage (semi-cured) and the second an additional curing stage (fully cured); the two series of vitrimers show complementary properties with excellent mechanical and thermal stability over a wide temperature range (250 °C), and they can each be effectively re-processed using thermal or chemical processing conditions. The produced set of vitrimers demonstrate an excellent combination of ultra-stretchability (up to 8000 for semi-cured and 3000 for fully cured vitrimers) and mechanical toughness (20–55 MJ/m3), where the balance between the two is tuneable by variation of the crosslink density and the processing conditions. This combination of stretchability and toughness significantly outperforms most other elastomers in the literature, and the semi-cured vitrimers also demonstrate good self-healing at room temperature. Moreover, the vitrimers show excellent mechanical damping properties with a maximum mechanical loss factor value of 2.0–3.0, which corresponds to the values for the best currently identified damping materials, and shows a significant loss factor 0.3 across a wide (and for application relevant) frequency range of 5 decades (0.001–100 Hz), or correspondingly across a wide -range (35 5 °C for 1 Hz) around room temperature. Both the processing conditions and the variation of the nominal cross-link density can be used to tune the -range of efficient mechanical loss. Table S3 (SI) provides a comparison between the properties of our vitrimers and those of high performance elastomers in literature.
Our vitrimers are promising candidates in polymer-based applications where stretchability combined with mechanical toughness and/or a high mechanical dissipation is required. Examples include applications in soft robotics or biomimetic synthetic materials, such as artificial skin, muscles, or cartilage, where a combination of high stretching and self-healing ability, together with a high mechanical strength and dissipation are required. The vitrimers are also suited to applications such as protective devices (helmets or clothing) or in vibration dampers or shock absorbers. Future work will focus on further developing the detailed understanding of the links between the primary polymer chain length and dynamic cross-link distributions and the resulting vitrimer physical properties, and to expand these investigations to the wider family of polyacrylate-based vitrimers.
@@ IV.4 Nuclear magnetic resonance (NMR) spectroscopy
1H NMR spectra were recorded on a Bruker AV4 NEO 11.75 T 500 MHz spectrometer fitted with a 5 mm Bruker C/H cryoprobe. For the 1H NMR experiments, the spectral width was 9090 Hz, the relaxation delay time was 2 s, and the number of scans was 16. Around 0.5 mg VMA and 0.7 ml DMSO- were first added to an NMR tube, followed by the addition of 5 l of D2O to ensure hydrolysis of the boronic ester crosslinker. The samples were left for 48h after sample preparation and then measured to ensure that the hydrolysis was complete. Chemical shifts (in ppm) are referenced to the DMSO- solvent peak with a chemical shift of 2.50 ppm.
@@ IV.5 Gel permeation chromatography (GPC)
The number average molecular weight (), weight average molecular weight (), and polymer dispersity () were measured using an Agilent 1260 Infinity system GPC instrument fitted with two 5 m Mixed-C columns plus a guard column, a refractive index detector, and a UV/Vis detector operating at 309 nm. THF (mixed with a small amount of water, vol/vol = 200:1) was used as the eluent (1 ml/min), and polystyrene standards ( ranging from 370 to 2,520,000 g/mol) were used for calibration. A polymer solution at a concentration of approximately 3 mg/ml was used for the measurements. The samples were left for 48h after sample preparation and then measured to ensure that hydrolysis was complete.
@@ IV.11 Time temperature superposition of SAOS rheology data
IV.12 Estimating the vitrimer mass density
The vitrimer samples were cut into well defined cuboids and weighed at room temperature (20 °C), which gave a mass density of 1.2 g/cm3. Given that the thickness of the vitrimer samples during SAOS experiments at 160 °C becomes ca 1.5 times that at 20 °C, and assuming that the sample cross-sectional areas do not change, the vitrimer densities at 160 °C were estimated to 0.8 g/cm3.
IV.13 Supporting Information
@@ IV.15 Conflict of Interest
IV.16 Data Availability Statement

### v2

@@ II.6 Mechanical damping
The -dependent loss tangent is plotted for all semicured (dashed lines) and fully cured (solid lines) vitrimer samples in Figure 5(h). A maximum loss factor () with a value between 2 and 3 is consistently observed. As the cross-link density increases upon full curing, the -relaxation slows down, as shown by the shift of the (Fig. 5g) and (Fig. 5h) loss peaks to higher temperatures, which is consistent with the increase in (see Table S2 in the SI). Also, increasing the cross-link density, either by addition of cross-linker during synthesis or by curing, shifts the peak in the loss tangent, and both curing and nominal cross-linking concentration can thus be used to control the -position of the mechanical loss maximum. The temperature span of efficient mechanical loss ( 0.3) cover 35 5 °C, which is comparable to the ranges ( °C) for commercial styrene-butadiene and natural rubbers ( 2.0). [75, 76] We note that the temperature range for efficient mechanical loss could be further increased by blending of semi-cured and fully cured vitrimers, thus producing materials with efficient loss factors across a range of 35 to 55 °C; such temperature ranges are comparable to those achieved for copolymer-based elastomers ( 1.7), [77] however, without the inherent compatibility issues that exist for blends of different rubbers ( 2.0). [78]
@@
II.7 Self-healing
@@
The photos in Figure 6(c) show a representative VMA-1’ vitrimer sample before deformation at 0% strain (left), after 1 hour at a strain of 2500 (centre), and after 24 hours at a strain of 3000 (right). In the left and centre photos, the cut can still be observed, but after 24 hours of room temperature self-healing it is no longer easily distinguishable, as shown in the right photo. Thus, we conclude that our semi-cured vitrimers demonstrate good self-healing properties at room temperature, without the addition of water or any other ‘solvent’. We do note that due to the slower segmental dynamics (ca. 20 °C higher ) and slower (circa two orders of magnitude; see Figure S8) network rearrangements, the fully cured samples do not exhibit noticeable room temperature self-healing on a time-scale of 24 hours. However, fully cured vitrimers have the potential to self-heal at temperatures above room temperature.
@@ III Conclusion
We have synthesized and characterised a new vitrimer system based on poly(methyl acrylate) with dynamic crosslinks utilising the associative exchange reactions of dioxaborolane metathesis. We demonstrate that these vitrimers combine ultra-stretchability and mechanical toughness with high mechanical dissipation and self-healing. Vitrimers with nominal cross-link densities between 2 and 8 mol% were prepared according to two different preparation routes: the first route involved a single curing stage (semi-cured) and the second an additional curing stage (fully cured); the two series of vitrimers showed complementary properties with excellent mechanical and thermal stability over a wide temperature range (250 °C), and they could each be effectively re-processed using thermal or chemical processing conditions. The produced set of vitrimers demonstrated an excellent combination of ultra-stretchability (up to 8000 for semi-cured and 3000 for fully cured vitrimers) and mechanical toughness (20–55 MJ/m3), where the balance between the two is tuneable by variation of the crosslink density and the processing conditions. This combination of stretchability and toughness significantly outperforms most other elastomers in the literature, and the lowest cross-link semi-cured vitrimers also demonstrate good self-healing at room temperature. Moreover, the vitrimers show excellent mechanical damping properties with a maximum mechanical loss factor value of 2.0–3.0, which corresponds to the values for the best currently identified damping materials, and shows a significant loss factor 0.3 across a wide (and for application relevant) frequency range of 5 decades (0.001–100 Hz), or correspondingly across a wide -range (35 5 °C for 1 Hz) around room temperature. Both the processing conditions and the variation of the nominal cross-link density can be used to tune the -range of efficient mechanical loss. Table S3 (SI) provides a comparison between the properties of our vitrimers and those of high performance elastomers in literature.
Our vitrimers are promising candidates in polymer-based applications where stretchability combined with mechanical toughness and/or a high mechanical dissipation is required. Examples include applications in soft robotics or biomimetic synthetic materials, such as artificial skin, muscles, or cartilage, where a combination of high stretching and self-healing ability, together with a high mechanical strength and dissipation are required. The vitrimers are also suited to applications such as protective devices (helmets or clothing) or in vibration dampers or shock absorbers. Future work will focus on further developing the detailed understanding of the links between the primary polymer chain length, the detailed dynamics and the distribution of cross-link, and the resulting vitrimer physical properties, as well as to expand these investigations to the wider family of polyacrylate-based vitrimers.
@@ IV.4 Nuclear magnetic resonance (NMR) spectroscopy
1H NMR spectra were recorded on a Bruker AV4 NEO 11.75 T 500 MHz spectrometer fitted with a 5 mm Bruker C/H cryoprobe. For the 1H NMR experiments, the spectral width was 9090 Hz, the relaxation delay time was 2 s, and the number of scans was 16. Around 0.5 mg VMA and 0.7 ml DMSO- were first added to an NMR tube, followed by the addition of 5 l of D2O to ensure hydrolysis of the boronic ester crosslinker. The samples were left for 48 hours after sample preparation and then measured to ensure that the hydrolysis was complete. Chemical shifts (in ppm) are referenced to the DMSO- solvent peak with a chemical shift of 2.50 ppm.
@@ IV.5 Gel permeation chromatography (GPC)
The number average molecular weight (), weight average molecular weight (), and polymer dispersity () were measured using an Agilent 1260 Infinity system GPC instrument fitted with two 5 m Mixed-C columns plus a guard column, a refractive index detector, and a UV/Vis detector operating at 309 nm. THF (mixed with a small amount of water, vol/vol = 200:1) was used as the eluent (1 ml/min), and polystyrene standards ( ranging from 370 to 2,520,000 g/mol) were used for calibration. A polymer solution at a concentration of approximately 3 mg/ml was used for the measurements. The samples were left for 48 hours after sample preparation and then measured to ensure that hydrolysis was complete.
@@ IV.11 Time temperature superposition of SAOS rheology data
IV.12 Supporting Information
@@ IV.14 Conflict of Interest
IV.15 Data Availability Statement